OpenStudy (anonymous):

what is the reciprocal of 4 2/5 22/5?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Very simple! To get the reciprocal of a number, just divide 1 by the number.

OpenStudy (anonymous):

69

OpenStudy (anonymous):

was i right?

OpenStudy (anonymous):

A 22/5 B 8/5 C 5/8 D 5/22

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

The reciprocal is found as follows: \[\large (4\small \frac{2}{5})\large ^{-1}=\frac{1}{4\small \frac{2}{5}}\large=\frac{1}{\frac{22}{5}}=\frac{5}{22}\]

OpenStudy (anonymous):

@kropot72 So I was wrong? Harmonyy just said I was right...

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kropot72):

@harmonyy is correct when saying: "To get the reciprocal of a number, just divide 1 by the number" but then contradicted that by confirming 22/5.

OpenStudy (anonymous):

just multiply it out (22/5); then flip it The answer is 5/22

OpenStudy (anonymous):

Distribute: 5(2x + 3) A. 10x + 15 B. 7x + 8 C. 10x + 25 D. 7x + 3